OpenStudy (anonymous):

solve by substitution: 2x+y=0 x-3y=-7 solve by elimination: 3x+4y=1 5x-3y=21

OpenStudy (anonymous):

Lets do the first one

OpenStudy (anonymous):

x-3y=-7 solve for x and tell me

OpenStudy (anonymous):

if you could take me through step by step that would be excellent, i have no idea how to do these at all, our teacher is out on maternity leave and she should have taught us this before she left

OpenStudy (anonymous):

yes that's what i am gonna do:)

OpenStudy (anonymous):

x-3y=-7 solve for x and tell me

OpenStudy (anonymous):

take the -3y on other side of the equation

OpenStudy (anonymous):

-4y

OpenStudy (anonymous):

Will it become x=-7+3y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I mean did u get the point? or shall i explain

OpenStudy (anonymous):

can you explain please?

OpenStudy (anonymous):

ok no problem

OpenStudy (anonymous):

x-3y=-7 is the equation right

OpenStudy (anonymous):

That is given in the question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

By adding +3y to both sides of the (=) we get x-3y+3y=-7+3y

OpenStudy (anonymous):

ok that makes sense

OpenStudy (anonymous):

Will -3y and +3y get cancelled out??

OpenStudy (anonymous):

i would say yes

OpenStudy (anonymous):

So, x=-7+3y

OpenStudy (anonymous):

Our second equation is 2x+y=0 which is given in the question

OpenStudy (anonymous):

What is the value of x we got?

OpenStudy (anonymous):

x=-7+3y

OpenStudy (anonymous):

So as the question says solve by substitution (put) Shall we put it in our new equation 2x+y=0 2(-7+3y)+y=0 <--- does that make sense?

OpenStudy (anonymous):

ok yes substitution makes sense.

OpenStudy (anonymous):

So, Multiply 2*-7 + 2*3y + y = 0

OpenStudy (anonymous):

so -14 +6y?

OpenStudy (anonymous):

-14 + 6y + y =0 -14 + 7y =0 7y = 14 Remember that signs change when we take somethin on the other side of the equation y = 14/7 y=2

OpenStudy (anonymous):

ok ok starting to make more sense!

OpenStudy (anonymous):

We got the value of y now we have to find the value of x

OpenStudy (anonymous):

tell me any of the equations from the question

OpenStudy (anonymous):

so we still have to figure out x for the first question?

OpenStudy (anonymous):

yes we got the value of y

OpenStudy (anonymous):

can we add -7+3y together?

OpenStudy (anonymous):

to make -4y?

OpenStudy (anonymous):

Well i don't mind teaching u but if u r completely new to this topic i would find it extremely difficult to explain u. I suggest it would be better if u ask some of your teachers or tutors because this topic ( Simultaneous equations) can't be taught through an online medium. Did u get my point.!